OpenStudy (anonymous):

find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. (5, 3) and (8, -4)

OpenStudy (anonymous):

\[slope = \frac{y_2-y_1}{x_2-x_1} = \frac{-4-3}{8-5} = \frac{-7}{3}\]